OpenStudy (anonymous):

Solve each equation with the quadratic formula: 10x^2-12x-2=0

OpenStudy (anonymous):

/ by 2 5x^2 - 6x -1 = 0 x =( --6 +- sqrt(36 - 4* 5 * -1) / 10

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

First imagine the formula as\[ax^2+bx+c=0\] then put it into the formula \[x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

I know the quadratic formula, and I forgot to divide everything by 2 in the beginning. Thanks, I can do it now.